use the method of diagonalization to obtain the general solution of the given system of differential equation d^2x1/dt^2 =5x1+4x2 and d^2x2/dt^2=x1+2x2


1
Expert's answer
2021-06-03T05:20:37-0400

Solution

Let vector X = (x1, x2)T , and matrix "A=\left(\begin{matrix}5&4\\1&2\\\end{matrix}\right)" .

Then system d2X/dt2 =AX.

det(A-λI) = (5-λ)(2-λ)-4 = λ2-7λ+6 = 0 =>  λ1=1, λ2=6

Matrix A may be reduced to diagonal matrix "D=\left(\begin{matrix}1&0\\0&6\\\end{matrix}\right)" by matrix S with columns, which are eigenvectors of systems (A - I)v1=0 and (A - 6I)v2=0.

Solving these systems we’ll get "S=\left(\begin{matrix}1&4\\-1&1\\\end{matrix}\right)"  and S-1AS=D

The system d2X/dt2 =AX is simplified by substitution Y=S-1X to the system d2Y/dt2 =DY with diagonal matrix D.

So "y_1=ae^t+be^{-t}", "y_2=ce^{t\sqrt{6}}+de^{-t\sqrt{6}}", where a,b,c,d – arbitrary constants.

X=SY => "x_1=ae^t+be^{-t}+4(ce^{t\sqrt{6}}+de^{-t\sqrt{6}})" , "x_2=-ae^t-be^{-t}+ce^{t\sqrt{6}}+de^{-t\sqrt{6}}"